OpenStudy (anonymous):

Giving medal An apple costs the same as 2 oranges. Together, an orange and a banana cost $0.10 more than an apple. Two oranges cost $0.15 more than a banana. What is the cost for an orange? a. $0.50 c. $0.25 b. $0.35 d. $0.10

OpenStudy (binary3i):

i ll tell you the equations you try solving it. A=2O O+B=.10 +A 2O= .15 +B

OpenStudy (anonymous):

b. O.o?

OpenStudy (binary3i):

substitute A in second equation so O + B = .10 + A O + B = .10 +2O B = .10 + O now add this equation to the last equation (B = .10 + O) + ( 2O= .15 + B) B + 2O = .10 + .15 + O + B O = .25

OpenStudy (anonymous):

ohhhh! ok thnx

OpenStudy (binary3i):

welcome.